1.Find the linear correlation coefficient for the paired data: (5,64), (10,86), (4,69), (6,86), (10,59), (9,87).

0.333

0.678

-0.224

0.750

0.999

-0.678

0.975

0.224 21.

2.Find the linear correlation coefficient for the paired data: (46.6,8), (27.2,8), (43.1,3), (29.0,9), (34.8,2).

0.2730

-0.181

0.543

0.307

-0.273

-0.307

-0.273

3. Find the equation of the regression line for the paired data: (6,2), (8,4), (20,13), (28,20), (36,30).

y = -3.79x + 0.897

y = -3.79 + 0.897

y = -2.79 + 0.897x

y = 0.897x - 3.79

y = -3.79 - 0.879x

y = -3.79x + 0.879

y = -3.79 - 0.897x

y = 3.79x + 0.879

4. Find the equation of the regression line for the paired data: (2,7), (4,11), (5,13), (6,20).

y = 0.15 + 2.8x

y = 0.15x + 3.0

y = 0.15x + 2.8

y = 0.15 + 3.0x

y = 3.0x y = 2.8x

y = 3.0 + x

y = 2.8 + x
Answer #1

1.x - Mx x-M )2 (Y - M (X- MxCY 124. 694 117.361 38.028 117.361 261.361 140.028 5 2.333 2.667 3.333 1.333 2.667 1.667 11.167 10.833 6.167 10.833 16.167 11.833 26.056 28.889 20.556 -14.444 -43.111 19.722 1.778 2.778 Sum: 35.333 Mx: 7.333 My: 75.167 Sum: 798.833 Sum: 37.667

X Values
∑ = 44
Mean = 7.333
∑(X - Mx)2 = SSx = 35.333

Y Values
∑ = 451
Mean = 75.167
∑(Y - My)2 = SSy = 798.833

X and Y Combined
N = 6
∑(X - Mx)(Y - My) = 37.667

R Calculation
r = ∑((X - My)(Y - Mx)) / √((SSx)(SSy))

r = 37.667 / √((35.333)(798.833)) = 0.2242